### v2

# Structuring Concept Space With The Musical Circle Of Fifths By Utilizing Music Grammar Based Activations

Abstract
In this paper, we explore the intriguing similarities between the structure of a discrete neural network, such as a spiking network, and the composition of a piano piece. While both involve nodes or notes that are activated sequentially or in parallel, the latter benefits from the rich body of music theory to guide meaningful combinations. We propose a novel approach that leverages musical grammar to regulate activations in a spiking neural network, allowing for the representation of symbols as attractors. By applying rules for chord progressions from music theory, we demonstrate how certain activations naturally follow others, akin to the concept of attraction. Furthermore, we introduce the concept of modulating keys to navigate different basins of attraction within the network. Ultimately, we show that the map of concepts in our model is structured by the musical circle of fifths, highlighting the potential for leveraging music theory principles in deep learning algorithms.
I Introduction
I-A Music Theory
I-A1 Background
Pythagoras, the ancient Greek mathematician and philosopher, made significant discoveries in the field of music that have had a lasting impact on our understanding of the relationship between mathematics and music. One of his most famous discoveries is the Pythagorean tuning system, which relates the frequencies of musical notes to simple ratios of whole numbers.
In the Pythagorean tuning system, the frequency of a note is determined by the ratio of the lengths of two vibrating strings. For example, if one string is twice as long as another, the note produced by the longer string will have a frequency that is half that of the note produced by the shorter string. This simple relationship between frequency and ratio forms the basis of the musical scale and allows for the creation of harmonious chords and melodies.
Pythagoras also discovered the concept of the musical circle of fifths, which is a diagram that shows the relationships between different musical keys. The circle of fifths is based on the mathematical concept of group theory, which is a branch of mathematics that studies the properties of groups of objects and the relationships between them.
In the circle of fifths, each key is represented by a point on the circle, and the keys are arranged in a way that shows the relationships between them based on the intervals of a perfect fifth. Moving clockwise around the circle, each key is a perfect fifth higher than the previous key, creating a harmonious progression of chords and melodies.
I-A2 Music Groups
Group theory is a branch of mathematics that deals with the study of groups, which are sets equipped with an operation that combines any two elements to produce a third element in the set. Groups have various properties, such as closure, associativity, identity element, and inverses, that make them a fundamental concept in mathematics.
In the context of music theory, group theory can be applied to analyze the structure and relationships between musical elements. One interesting application of group theory in music is the study of music groups, which are sets of musical elements (such as notes, chords, or scales) that form a group under a certain operation.
One example of a music group is the circle of fifths, which is a diagram that shows the relationships between the 12 tones of the Western musical scale. The circle of fifths is generated by the operation of moving up a perfect fifth from one note to the next, creating a cycle that eventually returns to the starting note.
The generators of the circle of fifths are the notes that form the basis of the cycle, such as the note C in the key of C major. By applying group theory concepts, we can analyze the structure of the circle of fifths and understand how the generators define the relationships between the different keys and chords in music.
Overall, group theory provides a powerful framework for studying the structure and relationships in music, and the analysis of music groups such as the circle of fifths can offer valuable insights into the underlying principles of music theory.
II Creating Musical Grammar Based Activations
II-A Introduction
We propose a novel approach that leverages musical grammar to regulate activations in a spiking neural network, allowing for the representation of symbols as attractors.
By applying rules for chord progressions from music theory, we show how certain activations naturally follow others, akin to the concept of attraction. Furthermore,we show the concept of modulating keys to navigate different basins of attraction within the network. Ultimately, the map of concepts in our model is structured by the musical circle of fifths.
Additionally, we propose a method to regularize a deep neural network by assigning a one-to-one mapping of the notes on a musical keyboard to the nodes on a neural network layer. This allows for the calculation of perceptual consonance in the activations of that layer.
We could train the neural network to provide an additional reward (proportional to the level of consonance in the layers)for agents. We further adapt the learning rule so that the value of perceptual consonance between nodes in different layers determines the learning rate of connections between them. Connections associated with higher perceptual consonance will have a higher learning rate than those with lower perceptual consonance.
The equations necessary are as follows:
Let $P_{ij}$ represent the perceptual consonance between nodes $i$ and $j$ in a neural network layer. The learning rate $\zeta_{ij}$ for the connection between nodes and $j$ is then determined by:
$\zeta_{ij}\propto P_{ij}$
When training a deep neural network we use an update rule for changing the weights as we learn. This can be interpreted as determining how much each weight $w_{jk}$ contributes to the error signal by weighting the error by the magnitude of the output activation from the previous (hidden) layer. The gradients with respect to each $w_{jk}$ are thus considered to be the “contribution” of that parameter to the total error signal and should be “negated” during learning. This gives the following gradient descent update rule for the output layer weights:
$w_{jk}$ $\leftarrow w_{jk}-\eta\frac{\partial E}{\partial w_{jk}}$ (1)
$\leftarrow w_{jk}-\eta({\delta_{k}}{a_{j}})$ (2)
Where $\eta$ is the learning rate E is the error of the network and $\delta_{k}$ can be interpreted as the network output error after being “back propagated” through the output activation function .In order to incorporate the learning rate $\zeta$ we need to add it to $\eta$.
We also believe that notes that are an octave apart, such as middle C and the C above it, share contexts of firing features and should be represented by related notes in the neural network. This could be or something that captures relationship.
II-B Concept Space
In this paper we take the view that populations of activations represent concepts. In fact a concept represents an attractor that modulates populations until it is stable. Just as an object falling into a basin. Here a population is a chord and basins are scales.
As we can see, with this system we are able to order the features detected by the neural network in a very organised and regularised manner.
Because D12, the music group , has 2 generators that span the group along different directions this makes a very complex feature representation system. Remember that within a layer, ALL the activations are related. And there are multiple ways in which they relate. This may be by them being a similar chord ,the same chord but different inversions, the same chord but higher extensions , the same chord but same scale, different chord but same chord type, chord in same scale , chordf rom parallel minor….and so forth.
Also the relationships between layers moves in the direction of the circle of fifths and fourths compounding relationships and regularizing activations across layers.
In fact since music theory is an unsolved theory, we have more relating activations than just the theory of music groups to ”file” concepts.
III Related Works
III-A Regularization
Regularization is a technique used in machine learning and deep learning to prevent overfitting and improve the generalization of a model. The earliest form of regularization is known as L1 regularization, also called Lasso regularization. L1 regularization adds a penalty term to the loss function that is proportional to the absolute value of the weights of the model. This encourages the model to learn sparse representations by setting some weights to zero.
The next form of regularization is L2 regularization, also known as Ridge regularization. L2 regularization adds a penalty term to the loss function that is proportional to the square of the weights of the model. This encourages the model to learn smooth representations by preventing the weights from becoming too large.
Another form of regularization is dropout, which randomly sets a fraction of the input units to zero during training. This helps prevent the model from relying too heavily on any one feature and encourages it to learn more robust representations.
Batch normalization is another regularization technique that normalizes the input to each layer of the neural network. This helps stabilize the training process and can improve the generalization of the model.
Orthogonal deep neural networks are a type of neural network architecture that enforces orthogonality constraints on the weight matrices of the network. This can help prevent the weights from becoming too correlated and can improve the stability and generalization of the model.
Overall, regularization techniques such as L1 and L2 regularization, dropout, batch normalization, and orthogonal deep neural networks can all help improve the generalization of deep learning models and prevent overfitting.
The techniques in this paper describe a method of regularization. In particular when surveying the literature our work is most closely related to that of orthogonal deep neural networks. Our work is similar but instead of separating features by a 90 degree angle we separate them by a 5th and a 4th according to the music group generators.
The advantages we gain are that the music group has 2 non trivial generators and we can have more complex dynamics as we pack representability into our networks.This should lead to greater generalisability and stability in the network.
III-B Grid cells and using the thousand brains theory
This work also pertains to the concept of grid cells in neuroscience. Grid cells are specialised neurons within animal brains that structure space and time for them by firing while the animal is moving. Certain grid cells fire after certain types of motion that cover space in a tesselation. With our ideas, we can view the circle of fifths and fourths as the grid that structures musical space, and as we move in it certain chords fire , signifying particular concepts or ideas, and positions on these two circles.
Grid cells are a type of neuron found in the entorhinal cortex of the brain that fire in a grid-like pattern as an animal moves through its environment. These cells play a crucial role in spatial navigation and are believed to be responsible for creating a cognitive map of the animal’s surroundings.
One of the leading theories in neuroscience is the Thousand Brains Theory proposed by Numenta, which suggests that the brain operates as a collection of thousands of mini-brains, each contributing to the overall perception and understanding of the world. In this theory, grid cells are seen as one of the key components that help the brain to create a unified representation of space and location.
Grid cells are thought to work in conjunction with other types of neurons, such as place cells and head direction cells, to form a comprehensive spatial representation in the brain. This network of cells allows animals to navigate their environment, remember locations, and plan future movements.
The relevance of grid cells to the Thousand Brains Theory lies in their ability to provide a consistent and accurate representation of space across different brain regions. By working together with other types of neurons, grid cells help to integrate information from various sensory inputs and cognitive processes, ultimately leading to a coherent understanding of the world.
In the context of deep learning methods, understanding the function of grid cells and their role in spatial navigation can inspire the development of more advanced artificial intelligence systems. By mimicking the principles of grid cells in neural networks, researchers can potentially improve the ability of AI systems to navigate complex environments, learn spatial relationships, and make more informed decisions based on their surroundings. This could have significant implications for a wide range of applications, from robotics to autonomous vehicles to virtual reality simulations.
In the context of the thousand brains theory,the implementation of grid cells is crucial for creating a fully functional biologically inspired autonomous system. Building upon the work of numenta, we propose the use of a spiking network with a columnar structure, where neurons are interconnected both within and between columns.
In our proposed network, the connections between neurons are not learned but designed to be highly symmetric across columns and sparse in connectivity. Each column in the network will output two vectors: one representing sensory information about an object and the other representing the object’s pose relative to the sense organ connected to that region of the network. Together, these vectors form the network’s hypothesis about the detected object.
To incorporate grid cells into the network, we will parse one copy of the input to the bottom layers and another copy will skip the first layer and be parsed to a middle layer. The neurons in this middle layer will represent the grid cells, with their output serving as the vector representing the object’s pose. The grid they will form will be structured by the circle of fifths and fourths.
To ensure that the network functions harmoniously, we will assign a mapping of neurons to a musical keyboard. This means that all the combined nodes in each layer of the spiking network, regardless of their column, should harmonize when mapped to musical notes. By doing so, we are essentially requiring that all the columns participate in representing the same musical structure, such as a chord, which in this context represents the detected object. Despite the harmonization requirement, the actual nodes firing between columns and within layers will be different, indicating that each column is forming a unique hypothesis about the object. This approach allows for the network to collectively detect and interpret objects while maintaining individual perspectives within each column.
Our method involves assigning each layer of a neural network its own three-dimensional space, where nodes represented by neurons move within the space as the network learns. The representation of a concept is then formed by a cluster of nodes within this space, which relates to the system’s disposition to form a particular chord during operation.
We also incorporate spiking graph neural networks with self-attention mechanisms.A spiking graph neural network with self-attention (SGNN) adapts its topology based on input.Each layer within each column will have a dedicated SGNN.The SGNN is trained to create a small graph with nodes positioned relative to their clusters within the three-dimensional space of the columns. Only those activations in the current time step will have representative clusters (at those angles ) within this SGNN so we will be attending to only those clusters that have been attracted.
The output of this graph serves as an object vector (combination of feature and pose vectors), which is compared to the output of a column ,selected by a voting mechanism ,as the final hypothesis. The column is then trained to output the correct hypothesis based on the error value.Taking the SGNN’s output to be the ” ground truth”.
The system dynamics involve initially unreliable hypotheses and activations in the columns, which are improved as the SGNN utilizes learned concepts to calculate the real object. As the system is connected to a global reward function, it learns to represent concepts more accurately over time and converges towards more faithful representations.
IV Conclusion
In conclusion, this paper, basing off of a population theory of concepts as attractors, shows how we can use a music grammar to represent these dynamics. chords then form populations, and scales handle the movement between chords or attraction. moving from one basin to the other as done in thought involves modulating keys and moving from one scale into another. This is more involved than has been described because music theory is as yet an unsolved theory. Later work will function on incorporating rhythm. This may be done with oscillatory spiking networks.More research needs to be done.We also discussed a method by which we could advance the work of numenta. By implementing the thousand brains theory from the perspective of music grammar activations and spiking grah neural networks with attention a full system may be realised.